The product of three positive consecutive integers is 15 times their sum find the three integers
Answer by CubeyThePenguin(3113)

(x-1)(x)(x+1) = 15((x-1) + x + (x+1))
x^3 - x = 15(3x)
x^3 - 46x = 0
x(x^2 - 46) = 0
x = 0, sqrt(46), or -sqrt(46)

The integers are -1, 0, and 1.